OpenStudy (gabbyalicorn):

Can someone check this? What is the value of y? 4y^2 + 8 = 264 I got 8.1 by doing 264/4 then to the square root but i feel like im suppose to do something is the ^2

Vocaloid (vocaloid):

almost

Vocaloid (vocaloid):

you need to isolate 4y^2 by subtracting 8 from both sides first

OpenStudy (gabbyalicorn):

oh so like 4y^2-8=264-8?

Vocaloid (vocaloid):

4y^2 + 8 - 8 = 264 - 8 4^2 = 256 y^2 = 64

Vocaloid (vocaloid):

(don't forget there's already a + 8 in the problem)

Vocaloid (vocaloid):

so y^2 = 64 y = +8 or - 8

OpenStudy (gabbyalicorn):

Ok thank you.